OpenStudy (anonymous):

The largest known prime number is \[\huge 2^{43112609} − 1\] the largest integer that is currently known to be a prime number. how many positive integers n are there for which \[\huge n^4+n^2+1\] is a prime

OpenStudy (shubhamsrg):

=> n^4 + 2n^2 -n^2 +1 = (n^2 +1)^2 -n^2 = (n^2 +n+1)(n^2 -n +1) since it is a factor of 2 terms, it wont be prime unless n^2 -n +1 = 1 or n^2 +n + 1 =1 =>n=0,1 and -1 since its asking for n>0, i'd say n=1 is the only ans..

OpenStudy (anonymous):

nice work @shubhamsrg

OpenStudy (shubhamsrg):

hmm.. :)